### v1

@@ 2.4 Counterfactual Explanations
[31] introduces a model‐agnostic method for generating counterfactual explanations in which the search is accelerated and interpretability is enhanced by guiding optimization with class prototypes. In their approach, prototypes are derived either as latent‐space centroids from an encoder or via class‐specific k‐d tree medoids, effectively replacing costly gradient‐based searches. The method was demonstrated on standard benchmarks, such as the MNIST image dataset and the Adult Census tabular dataset.
@@ 3 From SHAP-derived Rules to Sparse ECG Explanations
Our proposed method integrates rule-based feature selection with prototype-guided counterfactual generation to provide interpretable and clinically plausible explanations for ECG classification. As illustrated in Figure 1, the overall pipeline consists of three main components: (i) extraction of concise interval rules from SHAP importance values with Post-hoc Attribution Rule extraction [15], (ii) prototype selection via medoid clustering in DTW-based Multi-Dimensional Scaling (MDS) space, and (iii) an explanation module that aligns selected prototypes with the input signal and generates sparse counterfactuals. This modular structure enables fast and physiologically coherent explanations suitable for point-of-care use.
Figure 1: Overview of the proposed framework.
3.1 Extracting Rules from SHAP Values
We use Post-hoc Attribution Rule extraction (PHAR), a model-agnostic framework developed by [15] for converting numeric feature attributions into concise, instance-specific rules. PHAR supports numeric attributions from methods like SHAP [13] or LIME [19]. First, we select top features via a global threshold of the absolute attributions at a chosen percentile. Then we apply controlled perturbations within each feature’s variance, identifying stable value intervals that preserve the original prediction. The generated rules may facilitate transparent decision-making and support domain experts in validating model behavior, and can be visualized on time series as semi-factual scenarios [2]. The full PHAR pipeline, including rule fusion and optimization, is detailed in the original publication and is omitted here for brevity.
In this work, however, the rules from PHAR serve to determine the most important features in ECG signals and are subsequently used for clustering. In Section 3.2, these rules are used to generate prototypes by clustering the rule sets for each class and selecting medoids, where the ECG instance corresponding to a medoid rule becomes the prototype and serves as a counterfactual.
@@ 3.1.1 Conversion of Numeric Importance Values to Rules
We fix the global threshold $T$ at the 90.00th percentile of $E$:
@@
We then find, for each feature $f$, the smallest interval $(l_{f},h_{f}]$ such that every perturbed sample produces the same full prediction vector $\mathbf{y}_{n}$. This enforces consistency across all output labels.
@@ 3.2 Prototype Selection
    This filters samples that belong exclusively to class $l$. This step excludes ambiguous or potentially misclassified recordings, yielding prototypes that are more stable, representative, and clinically meaningful.
@@
    Distance-based Dimensionality Reduction: In this step, we compute pairwise Dynamic Time Warping (DTW) distances between all training samples within each class, creating a distance matrix that captures morphological similarities between ECG signals.
    For two ECG signals $x_{i},x_{j}\in\mathcal{D}_{l}$, the Dynamic Time Warping distance is:
[…]
@@
    where $a(i)$ is the mean intra-cluster distance and $b(i)$ is the mean nearest-cluster distance.
@@
    Clustering in MDS Space: We partition $Z_{l}$ into $k$ clusters using k-means:
@@ 3.3 From Prototypes to Sparse ECG Explanations
Given a query ECG signal and its model prediction, our method generates counterfactual explanations through a two-stage process: prototype alignment followed by sparsification. This approach ensures both temporal coherence and minimal perturbations.
@@ 3.3.1 R-peak Based Temporal Alignment
ECG signals exhibit natural variability in heart rate and rhythm, even within the same diagnostic class. To generate physiologically plausible counterfactuals that can be more easily analyzed by overlaying them, we align the temporal structure of selected prototypes to match the rhythm of the query signal. We employ R-peak detection and piecewise-linear warping:
@@
    • When the prototype has fewer beats, we prepend zeros to its start and, in the alignment step, only warp the beats that correspond to those in the query signal.
    • When the prototype has more beats, we trim it to match the query signal’s beat count.
@@
$\mathbf{x}_{\text{aligned}}[t_{k}:t_{k+1}]=\text{interp1d}(\mathbf{x}_{\text{proto}}[s_{k}:s_{k+1}],[0,1]\rightarrow[0,1])$ (12)
where $t_{k},t_{k+1}$ are consecutive R-peak positions in the query, and $s_{k},s_{k+1}$ are the corresponding peaks in the prototype.
@@ 3.3.2 Sparsity Optimization
The aligned prototype serves as an initial counterfactual, but it differs from the query at every time point. To enhance interpretability and identify the minimal changes required for class transition, we apply sparsity optimization:
@@
    Importance scoring: We compute the absolute difference between the aligned prototype and query at each time-channel point. R-peak regions receive doubled importance scores to preserve critical cardiac features.
@@
where $\mathbf{m}$ is the binary modification mask, $\mathbf{x}_{q}$ is the query, $\mathbf{x}_{p}$ is the aligned prototype, $\odot$ denotes element-wise multiplication, and $f$ is the model.
@@ 3.3.3 Query Processing Workflow
    Retrieve the pre-computed prototypes for the target class, which were extracted during the offline MDS-based clustering phase. Select the prototype with minimal DTW distance to the query to choose a prototype that shares similar morphological characteristics with the input signal.
@@
This approach generates explanations in under a second by using pre-computed prototypes. The two-stage process, temporal alignment followed by sparse modifications, produces counterfactuals that respect ECG physiology. By preparing prototypes offline and running only lightweight optimizations at query time, we achieve both speed and quality, making the system practical for point-of-care use.
@@ 3.4 Objective Evaluation
To comprehensively assess the quality of generated counterfactual explanations, we developed five key evaluation metrics for 12-lead ECG explanations, which are commonly used in the literature for tabular counterfactuals ([5]), which we adapted for the specific characteristics of multivariate timeseries ECG signals. Unlike tabular data, where features are independent, ECG signals exhibit temporal dependencies, morphological constraints, and channel correlations that require specialized evaluation approaches.
@@ 3.4.1 Validity
Validity is the fundamental requirement for any counterfactual explanation, measuring whether it successfully changes the model’s prediction to the desired target class. For ECG signals, we must ensure that the counterfactual not only crosses the decision boundary but also maintains physiological plausibility. Given a query ECG signal $\mathbf{x}\in\mathbb{R}^{T\times C}$ (where $T=1000$ samples and $C=12$ leads) with model prediction $f(\mathbf{x})=y$, and its counterfactual $\mathbf{x}^{\prime}$, validity is defined as:
$\text{validity}(\mathbf{x},\mathbf{x}^{\prime})=\begin{cases}1&\text{if }f(\mathbf{x}^{\prime})=y_{\text{target}}\land f(\mathbf{x}^{\prime})\neq f(\mathbf{x})\\ 0&\text{otherwise}\end{cases}$ (14)
@@
$\text{validity}_{\text{multi}}=\prod_{i=1}^{L}\mathbb{1}[f_{i}(\mathbf{x}^{\prime})=y_{\text{target},i}]$ (15)
where $L$ is the number of diagnostic labels and $f_{i}$ represents the $i$-th class prediction.
@@ 3.4.2 Sparsity
Sparsity quantifies the locality and efficiency of modifications, crucial for interpretability in clinical settings ([4]). Sparse counterfactuals highlight specific ECG features (e.g., ST-segment changes, QRS morphology) responsible for diagnostic decisions. We evaluate sparsity through multiple complementary metrics:
@@
    where the threshold $\tau=0.01\cdot\sigma_{\mathbf{x}}$ accounts for noise tolerance, with $\sigma_{\mathbf{x}}$ being the standard deviation of the training data.
@@ 3.4.3 Stability
Stability metrics assess the robustness of counterfactual explanations against minor perturbations ([6]), crucial for clinical reliability where measurement noise and physiological variations are inherent.
@@
    $\text{stability\_score}=\frac{1}{|\mathcal{N}|\cdot N_{\text{trials}}}\sum_{i=1}^{N_{\text{trials}}}\sum_{\sigma_{\epsilon}\in\mathcal{N}}\mathbb{1}[f(\mathbf{x}^{\prime}+\epsilon_{i})=f(\mathbf{x}^{\prime})]$ (20)

### v2

@@ 2.5 Prototype-based Explanations
Prototypes, serving as class representatives to guide the search for counterfactual explanations, were discussed by [vanlooveren2020interpretablecounterfactualexplanationsguided]. They introduced a model-agnostic method for generating counterfactual explanations in which the search is accelerated and interpretability is enhanced by guiding optimization with class prototypes. In their approach, prototypes are derived either as latent-space centroids from an encoder or via class-specific k-d tree medoids, effectively replacing costly gradient-based searches. While their work focused on static benchmarks like MNIST or the Adult Census dataset, it established the foundation for using exemplars to improve both speed and interpretability.
Extending the concept of prototype guidance to the temporal domain, [Li2025Reliable] recently proposed a framework utilizing ShapeDBA. This technique employs Dynamic Time Warping (DTW) on local structural features to compute a representative average (barycenter) that minimizes temporal distortion. However, their validation relied on the TwoLeadECG dataset from the UCR Archive, which typically consists of short, pre-segmented heartbeats. In contrast, clinical 12-lead recordings involve continuous signals with variable rhythms and multiple cardiac cycles. While ShapeDBA mitigates alignment issues in short segments, our framework employs real patient medoids and explicit R-peak alignment to strictly preserve physiological fidelity across complex recordings, avoiding the potential smoothing artifacts of barycentric averaging.
Furthermore, effective prototype-based reasoning requires not just identifying a similar example, but explaining why it is relevant.  [Karolczak2025] argue that users must understand which specific features make a prototype similar to the query instance. They introduced the concept of “alike parts”, which are segments identified by integrating feature importance scores from model-agnostic methods directly into the prototype selection process. By demonstrating that fusing such importance scores, derived from various attribution techniques like SHAP or LIME, with example-based explanations significantly enhances user understanding and perceived transparency, their work provides a strong rationale for our approach. This validates our strategy of utilizing importance-driven highlights to identify specific ECG regions, thereby streamlining the clinical reasoning process and ensuring that diagnostic waveform alterations are immediately apparent to the practitioner.
Synthesizing these perspectives, the literature establishes that seamless workflow integration and interactive visualization are essential for translating the high performance of deep learning models into clinical practice. However, a critical gap persists in generating explanations that are both physiologically plausible and rapidly interpretable. While conventional counterfactual methods often sacrifice biological realism to minimize the numerical distance to the query instance, and existing prototype-based approaches lack the precise temporal alignment required for complex ECG rhythms, our framework addresses these limitations. By combining real-patient prototypes with explicit R-peak synchronization and sparsity optimization, we propose a solution that is physiologically consistent, visually intuitive, and designed to facilitate clinical integration and support diagnostic workflows.
@@ 3 From SHAP-derived Rules to Sparse ECG Explanations
Our proposed method integrates rule-based feature selection with prototype-guided counterfactual generation to provide interpretable and clinically plausible explanations for ECG classification. As illustrated in Figure 1, the overall pipeline consists of three main components: (i) extraction of concise interval rules from SHAP importance values with Post-hoc Attribution Rule extraction [mozolewski2025phar], (ii) prototype selection via medoid clustering in DTW-based Multi-Dimensional Scaling (MDS) space, and (iii) an explanation module that aligns selected prototypes with the input signal and generates sparse counterfactuals. This modular structure enables fast and physiologically coherent explanations suitable for point-of-care use.
3.1 PHAR: Rule Extraction from Feature Attributions
We use Post-hoc Attribution Rule extraction (PHAR), a model-agnostic framework developed by [mozolewski2025phar] for converting numeric feature attributions into concise, instance-specific rules. PHAR supports numeric attributions from methods like SHAP [Lundberg2017] or LIME [LIME]. These attributions serve only as a preliminary filter to identify regions of interest for PHAR. First, we select top features via a global threshold of the absolute attributions at a chosen percentile. Then we apply controlled perturbations within each feature’s variance, identifying stable value intervals that preserve the original prediction. While the framework is explainer-agnostic, we adopt SHAP as the importance source due to its solid game-theoretic foundations. We prioritize this approach because it ensures that the initial attribution is theoretically distinct from the perturbation-based logic used in PHAR’s subsequent interval-identification step. Any potential noise in the underlying explainer is mitigated by PHAR’s internal consistency checks, which require prediction invariance within the identified intervals. Ultimately, the transition from abstract rule-based importance to explanations anchored in real-patient medoids (Section 3.2) ensures that the final insights are not only stable but also physiologically plausible.
The generated rules may facilitate transparent decision-making and support domain experts in validating model behavior, and can be visualized on time series as semi-factual scenarios [semi_ijcai2023p732]. In this work, however, the rules from PHAR serve to determine the most important features in ECG signals and are subsequently used for clustering. In Section 3.2, these rules are used to generate prototypes by clustering the rule sets for each class and selecting medoids, where the ECG instance corresponding to a medoid rule becomes the prototype and serves as a counterfactual. To ensure this paper is self-contained, Section 3.1.1 provides a comprehensive overview of the core PHAR mechanism as adapted for our ECG classification framework, while further details on rule fusion strategies can be found in [mozolewski2025phar].
@@ 3.1.1 Conversion of Numeric Importance Values to Rules
The global threshold $T$ is set to the 90.00th percentile of the distribution $E$.
@@
This choice is grounded in our ablation study (see Appendix A.3), which demonstrates that the 90th percentile yields the highest pattern granularity while optimizing computational time, avoiding the loss of informative signal observed at higher extremes.
@@
We then find, for each feature $f$, the smallest interval $(l_{f},h_{f}]$ such that every perturbed sample produces the same full prediction vector $\mathbf{y}_{n}$. This enforces consistency across all output labels. In rare cases where no stable interval exists (i.e., all perturbations alter the prediction), the feature is excluded from the rule, as it indicates a highly sensitive decision boundary region where small changes cause class transitions.
@@
Here, $R_{n}$ denotes the rule for instance $n$, $\bigwedge$ represents the logical conjunction over the set of important features $F_{*}$. For each important feature $f$, the value $x_{f}$ must fall within ($l_{f}$, $h_{f}$] to preserve the original prediction.
@@ 3.2 Prototype Selection
    Here, $\mathcal{D}_{l}$ denotes the filtered subset for class $l$, $(x_{i},y_{i})$ represents a sample-label pair from the training set $\mathcal{D}_{\text{train}}$, $y_{i,l}$ is the binary label for class $l$, and the summation condition $\sum_{j=1}^{L}y_{i,j}=1$ enforces that the selected sample is single-label. This filters samples that belong exclusively to class $l$. This design choice prioritizes generating clear, actionable counterfactuals that highlight minimal changes for a single diagnostic transition, aligning with how clinicians typically reason about differential diagnosis. The comorbidity combinations are common (Table 2), but treating each combination as a separate target class would exponentially increase prototype complexity and reduce interpretability. Furthermore, filtering for single-class exclusivity implicitly removes ambiguous or potentially noisy recordings, ensuring that the resulting prototypes are stable and highly representative of their core pathology.
@@
    Distance-based Dimensionality Reduction: In this step, we compute pairwise Dynamic Time Warping (DTW) distances between all training samples within each class, creating a distance matrix that captures morphological similarities between ECG signals. DTW was selected for its ability to accommodate natural heart rate variability, aligning corresponding waveform components even across patients with different cardiac rhythms. For two ECG signals $x_{i},x_{j}\in\mathcal{D}_{l}$, the Dynamic Time Warping distance is:
    $\text{DTW}(x_{i},x_{j})=\min_{\pi\in\Pi}\sum_{(t,s)\in\pi}\|x_{i}(t)-x_{j}(s)\|_{2}$ (7)
@@
    Where $z_{i}$ is the d-dimensional embedding of ECG sample $x_{i}$, and $Z_{l}$ is the matrix of all embeddings for class $l$, and $a(i)$ is the mean intra-cluster distance and $b(i)$ is the mean nearest-cluster distance.
@@
    Clustering in Multi-Dimensional Scaling (MDS) Space: We partition $Z_{l}$ into $k$ clusters using k-means (the empirical selection of the hyperparameter $k$ via silhouette analysis is detailed in Section 4.7.3):
@@
    By selecting medoids (actual training instances) rather than abstract centroids, we ensure that every explanation is anchored in a real, verifiable patient record. This provides a traceable data origin essential for governance in safety-critical clinical systems, distinguishing our approach from generative methods that may synthesize plausible yet unverified signal patterns.
@@ 3.3 From Prototypes to Sparse ECG Explanations
Given a query ECG signal and its model prediction, our method generates counterfactual explanations through a two-stage process: prototype alignment followed by sparsification. This approach ensures both temporal coherence and minimal perturbations. To ensure physiological plausibility, we implement a specialized segment cleaning procedure detailed in Section 3.3.2, which enforces a minimum length of $10$ samples to eliminate scattered or isolated modifications that lack clinical meaning. Furthermore, by generating modifications across consecutive timestamps and anchoring the explanations in real-patient medoids, we ensure that the final insights are both stable and grounded in empirical cardiac morphologies rather than stochastic artifacts.
@@ 3.3.1 R-peak Based Temporal Alignment
ECG signals exhibit natural variability in heart rate and rhythm, even within the same diagnostic class. Consequently, a direct element-wise comparison between a query and a prototype is often erroneous due to phase shifts or mismatches in the number of beats. To address this, we adopt a query-centric approach: since the explanation must interpret the specific patient’s recording, we strictly maintain the temporal structure of the query signal. To generate physiologically plausible counterfactuals that can be more easily analyzed by overlaying them, we align the temporal structure of selected prototypes to match the R-peaks of the query signal. The alignment procedure warps the prototype’s time axis to match the query’s R-peak positions, ensuring that morphological comparisons focus on shape differences within corresponding cardiac cycles rather than rhythm discrepancies. We employ R-peak detection and piecewise-linear warping:
@@
    • When the prototype has fewer beats, we pad the regions preceding the first and following the last matching R-peak with zeros to match the query duration. In the subsequent alignment step, we only warp the beats that correspond to those in the query signal, leaving the padded boundary regions flat. Although this introduces local non-physiological artifacts, it ensures that the complete diagnostic context of the query is preserved. Crucially, these artifacts are typically discarded during the sparsity optimization (Section 3.3.2), as they do not contribute to the target classification.
    • When the prototype has more beats, we truncate the preceding cardiac cycles to match the query’s beat count. However, as the piecewise-linear warping is mathematically defined strictly between corresponding R-peaks, the signal regions outside this interval, specifically before the first and after the last matched peak—are set to zero. These resulting flat boundaries are unavoidable in peak-to-peak alignment but are subsequently discarded by the sparsity optimization, ensuring the final explanation remains physiologically plausible.
@@
$\mathbf{x}_{\text{aligned}}[t]=\mathbf{x}_{\text{proto}}\left(s_{k}+\frac{s_{k+1}-s_{k}}{t_{k+1}-t_{k}}(t-t_{k})\right)$ (12)
where $t_{k},t_{k+1}$ are consecutive R-peak positions in the query, and $s_{k},s_{k+1}$ are the corresponding peaks in the prototype. Since the mapped time coordinate is continuous, the value of $\mathbf{x}_{\text{proto}}$ is evaluated via linear interpolation. This linear scaling effectively stretches or compresses the prototype segment to perfectly match the duration of the query’s cardiac cycle.
@@ 3.3.2 Sparsity Optimization
The aligned prototype serves as an initial counterfactual, but since it originates from a different patient recording, it inherently differs from the query at nearly every time point despite temporal alignment. To enhance interpretability and identify the minimal changes required for class transition, we apply sparsity optimization:
@@
    Importance scoring: We compute the absolute difference between the aligned prototype and query at each time-channel point. Crucially, R-peak regions receive doubled importance scores to prioritize central cardiac features over peripheral signal components. Consequently, the flat boundary artifacts introduced by alignment (Section 3.3.1) typically yield lower weighted scores compared to high-magnitude diagnostic discrepancies (e.g., ST-segment deviations), causing them to be naturally deprioritized during selection.
@@
where $\mathbf{m}$ is the binary modification mask, $\mathbf{x}_{q}$ is the query, $\mathbf{x}_{p}$ is the aligned prototype, $\odot$ denotes element-wise multiplication, and $\mathcal{M}$ is the model.
@@ 3.3.3 Query Processing Workflow
    Retrieve the pre-computed prototypes for the target class, which were extracted during the offline MDS-based clustering phase. Select the prototype with minimal DTW distance to the query. As discussed in Section 2, we prioritize existing patient medoids over synthetic averages (barycenters) to ensure that the baseline explanation remains physiologically authentic and retains all high-frequency clinical details.
@@
This approach generates explanations in under a second per 10-second ECG recording (1000 samples at 100Hz) by using pre-computed prototypes. The two-stage process, temporal alignment followed by sparse modifications, produces counterfactuals that respect ECG physiology. By preparing prototypes offline and running only lightweight optimizations at query time, we achieve both speed and quality. This design specifically targets the low-latency requirements of real-time monitoring systems, achieving a balance between physiological fidelity and the rapid response times necessary for point-of-care deployment.
@@ 3.4 Objective Evaluation
To comprehensively assess the quality of generated counterfactual explanations, we developed five key evaluation metrics for 12-lead ECG explanations, which are commonly used in the literature for tabular counterfactuals ([bayrak2024evaluation]), which we adapted for the specific characteristics of multivariate timeseries ECG signals. Unlike tabular data, where features are independent, ECG signals exhibit temporal dependencies, morphological constraints, and channel correlations that require specialized evaluation approaches.
@@ 3.4.1 Validity
Validity is the fundamental requirement for any counterfactual explanation, measuring whether it successfully changes the model’s prediction to the desired target class. For ECG signals, we must ensure that the counterfactual not only crosses the decision boundary but also maintains physiological plausibility. Given a query ECG signal $\mathbf{x}\in\mathbb{R}^{T\times C}$ (where $T=1000$ samples and $C=12$ leads) with model prediction $\mathcal{M}(\mathbf{x})=y$, and its counterfactual $\mathbf{x}^{\prime}$, validity is defined as:
$\text{validity}(\mathbf{x},\mathbf{x}^{\prime})=\begin{cases}1&\text{if }\mathcal{M}(\mathbf{x}^{\prime})=y_{\text{target}}\land\mathcal{M}(\mathbf{x}^{\prime})\neq\mathcal{M}(\mathbf{x})\\ 0&\text{otherwise}\end{cases}$ (14)
@@
$\text{validity}_{\text{multi}}=\prod_{i=1}^{L}\mathbb{1}[\mathcal{M}_{i}(\mathbf{x}^{\prime})=y_{\text{target},i}]$ (15)
where $L$ is the number of diagnostic labels, $\prod$ denotes the product over all classes (equivalent to a logical AND), $\mathbb{1}[\cdot]$ is the indicator function returning 1 if the condition holds and 0 otherwise, $\mathcal{M}_{i}(\mathbf{x}^{\prime})$ represents the model’s binary prediction for class $i$ given counterfactual $\mathbf{x}^{\prime}$, and $y_{\text{target},i}$ is the desired label for that class.
@@ 3.4.2 Sparsity
Sparsity quantifies the locality and efficiency of modifications, crucial for interpretability in clinical settings ([bayrak2023pertcf]). Sparse counterfactuals highlight specific ECG features (e.g., ST-segment changes, QRS morphology) responsible for diagnostic decisions. We evaluate sparsity through multiple complementary metrics:
@@
    where $T$ and $C$ denote the number of time points and leads respectively, $\mathbf{x}^{\prime}_{t,c}$ and $\mathbf{x}_{t,c}$ are the signal values at time $t$ and lead $c$ for the counterfactual and query, and $\tau=0.01\cdot\sigma_{\mathbf{x}}$ is the noise tolerance threshold derived from the training data standard deviation $\sigma_{\mathbf{x}}$.
@@ 3.4.3 Stability
Stability metrics assess the robustness of counterfactual explanations against minor perturbations ([darias2024empirical]), crucial for clinical reliability where measurement noise and physiological variations are inherent.
@@
    $\text{stability\_score}=\frac{1}{|\mathcal{N}|\cdot N_{\text{trials}}}\sum_{i=1}^{N_{\text{trials}}}\sum_{\sigma_{\epsilon}\in\mathcal{N}}\mathbb{1}[\mathcal{M}(\mathbf{x}^{\prime}+\epsilon_{i})=\mathcal{M}(\mathbf{x}^{\prime})]$ (20)